OpenStudy (anonymous):

(5x^3 y^6) (6x^2 y^5) simplify how can I do this showing steps to get answer?

OpenStudy (anonymous):

To do this equation we must distribute the whole left parentheses. (5x^3y^6)(6x^2y^5) Remember when multiply exponents with the same variable, we add them. 30x^5y^11. I got that by multiplying each common term. 5x6=30 We have our exponents 3 and 2, which give us the exponent of 5. Since y is not a coefficient we just add our 2 exponents. Our exponents are 6 and 5, which gives us the exponent of 11. FINAL ANSWER: 30x^5y^11.

hero (hero):

\[\color {green} {\huge{\checkmark}}\]

OpenStudy (anonymous):

Your welcome! :D

OpenStudy (anonymous):

roseanne! hello

OpenStudy (anonymous):

thank you very much!

hero (hero):

I checked the answer, so it's good

OpenStudy (anonymous):

Hey hero I have a final tom orrow for algebra could u help me a lil bit?

OpenStudy (anonymous):

Then give me a medal. :3

hero (hero):

um...um...

hero (hero):

What time?

OpenStudy (anonymous):

is my final?

hero (hero):

What time are you doing it, yes.

OpenStudy (anonymous):

11:30am

OpenStudy (anonymous):

what time zone?

OpenStudy (anonymous):

Florida

OpenStudy (anonymous):

florida daylight time?

OpenStudy (anonymous):

sat will you help me show me steps on these problems

OpenStudy (anonymous):

name one

OpenStudy (anonymous):

simplify (w^19) ^0 (w^6)

hero (hero):

w^6

OpenStudy (anonymous):

anything to the power of zero is 1. so ignore that one

OpenStudy (anonymous):

you are left with just what hero said, \[w^6\]

OpenStudy (anonymous):

It always will equal to 1 when the exponent is 0. Regardless of the numbers.

OpenStudy (anonymous):

whenever you see \[b^0\] remember that \[b^0=1\]

OpenStudy (anonymous):

so how is it w^6?

OpenStudy (anonymous):

Regardless of terms, always will be 1. 6\[6x ^{0} \]= 1

OpenStudy (anonymous):

we go slow. \[(w^{19})^0=1\]

OpenStudy (anonymous):

It is w^6 because w^6 x 1= w^6. XD

OpenStudy (anonymous):

Any positive rational number times itself will always be the same.

OpenStudy (anonymous):

so \[(w^{19})^0\times w^6=1\times w^6=w^6\]

OpenStudy (anonymous):

(Y) Good job! :3 Anymore quesions?

OpenStudy (anonymous):

ok multiply (4x+5) (2x^2-x-3)

OpenStudy (anonymous):

first you have to multiply everything in the second parentheses by 4x \[4x(2x^2-x-3)\]

OpenStudy (anonymous):

when you do this, multiply the numbers together, and add one to each exponent. you will get \[8x^3-2x^2-12x\] is that part clear? we are not done yet

OpenStudy (anonymous):

To multiply it we distribute each term in the left parentheses to the right. 4x2=8 Multiply exponents: 1x2 or 1+2=3 8x^3-4x-12x Above is the first term distributed to the whole group. Now the 2nd term. 10x^2-5x-15. Combine like terms. 8x^3-4x-12+10x^2-5x-15 8x^3-9x-27 is our final answer.

OpenStudy (anonymous):

then you have to multiply everything in the second parentheses by 5. you get \[5(2x^2-x-3)\] which is \[10x^2-5x-15\]

OpenStudy (anonymous):

so in total you have \[8x^3-2x^2-12x+10x^2-5x-15\]

OpenStudy (anonymous):

now you need to combine like terms. \[8x^3\] has no other like terms because the exponent is 3 and it is the only term with an exponent of 3

OpenStudy (anonymous):

\[-2x^2+10x^2=8x^2\] because they are like terms and \[-2+10=8\]

OpenStudy (anonymous):

\[-12x-5x=-17x\] for the same reason, and -15 has no like term. so your "final answer" is \[8x^3+8x^2-17x-15\]

OpenStudy (anonymous):

except that i messed up on the x squared term. hold on

OpenStudy (anonymous):

graph the line which is easy but how do I convert y=2x+3 to a point?

OpenStudy (anonymous):

hold on one second i made a mistake \[4x(2x^2-x-3)\] should be \[8x^3-4x^2-12x\] so your answer above should be \[8x^3+6x^2-17x-15\]

OpenStudy (anonymous):

you do not convert \[y=2x+3\] so a point. but if you want a point on the line you pick any x, and then find y

OpenStudy (anonymous):

for example if i pick x = 1 then i know \[y=2\times 1+3=2+3=5\] so the point (1,5) is on the line

OpenStudy (anonymous):

To convert it to a point: we use our y intercept and slope. Let's start with y intercept which is (0,3). Then use slope (rise/run.) Our rise is 2/1 From our y intercept we move up the y axis 2 point and to the right one point, which is point (1,5). Then connect the lines.

OpenStudy (anonymous):

if i pick x = 2 then i know \[y=2\times 2+3=4+3=7\] so the point (2,7) is on the line. and so on

OpenStudy (anonymous):

subtract (10x^2+7x-2) - (4x^2-2x+2)